# Continuous-Time Value Iteration for Multi-Agent Reinforcement Learning

### v2

@@ Appendix A Mathematical Proof
Given all $(x,t)\in\mathcal{X}\times[0,T]$ and a small horizon $\Delta t>0$, we apply a Taylor expansion to the definition of $V(x,t)$ in Eq. (2) to derive the HJB PDE as follows:
$V(x,t)$ $=\max_{u\in\mathcal{U}}\int_{t}^{T}e^{-\rho(\tau-t)}r(x,u)d\tau+g(x_{T})$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+g(x_{T})$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t}^{T-\Delta t}e^{-\rho(s+\Delta t-t)}r(x(s+\Delta t),u(s+\Delta t))ds+g(x_{T})$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}(\max_{u\in\mathcal{U}}\int_{t}^{T-\Delta t}e^{-\rho(s-t)}r(x(s+\Delta t),u(s+\Delta t))ds+g(x_{T}))$
$\approx\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}(\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{T}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+g(x_{T}))$
$\approx\max_{u\in\mathcal{U}}r(x,u)\Delta t+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}V(x(t+\Delta t),t+\Delta t)$
$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+(1-\rho\Delta t+o(\Delta t))\max_{u\in\mathcal{U}}(V(x,t)+\partial_{t}V(x,t)\Delta t+\nabla_{x}V(x,t)^{\!\top}\!f(x,u)\Delta t+o(\Delta t))$
By canceling out $V(x,t)$ on both sides of the above equality, we obtain that
$-\rho V(x,t)\Delta t+\partial_{t}V(x,t)\Delta t+\max_{u\in\mathcal{U}}(\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u))\Delta t+o(\Delta t)=0$
@@
$-\rho V(x,t)+\partial_{t}V(x,t)+\max_{u\in\mathcal{U}}(\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u))=0.$
Therefore, the $V(x,t)$ is the optimal solution to the following HJB PDEs:
$\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f(x,u^{*})+r(x,u^{*})=0,$
with boundary condition $V(x,T)=g(x)$.
Here optimal control input is $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,t,\nabla_{x}V(x,t))$, where $\mathcal{H}$ is the Hamiltonian defined as $\mathcal{H}=\nabla_{x}V(x,t)^{\!\top}\!f(x,u)+r(x,u)$. ∎
@@
$Q(x,t,u)=r(x,u)\,\delta t+e^{-\rho\delta t}\,V\bigl(x(t+\delta t),\,t+\delta t\bigr).$
For small $\delta t$ we have the first-order Taylor expansion in both state and time:
$V\bigl(x(t+\delta t),\,t+\delta t\bigr)=V(x,t)+\nabla_{x}V(x,t)^{\!\top}\!f(x,u)\,\delta t+\partial_{t}V(x,t)\,\delta t+o(\delta t).$
@@
$Q(x,t,u)$ $=r(x,u)\,\delta t+\bigl(1-\rho\,\delta t+o(\delta t)\bigr)\Bigl[V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)\,\delta t+\partial_{t}V(x,t)\,\delta t+o(\delta t)\Bigr]$
$=r(x,u)\,\delta t+V(x,t)+\bigl[\nabla_{x}V(x,t)^{\!\top}f(x,u)+\partial_{t}V(x,t)-\rho V(x,t)\bigr]\delta t+o(\delta t).$
Subtract $V(x,t)$ and discard the higher-order term:
$Q(x,t,u)-V(x,t)=[-\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)+r(x,u)]\delta t+o(\delta t).$
@@
$A(x,t,u)=\lim_{\delta t\to 0}\frac{Q(x,t,u)-V(x,t)}{\delta t}=-\rho V(x,t)+\partial_{t}V(x,t)+\nabla_{x}V(x,t)^{\!\top}f(x,u)+r(x,u).$
@@
$\pi_{\text{new}}(x,t)=\arg\max_{u\in\mathcal{U}}Q^{\pi_{\text{old}}}(x,t,u),$
@@
$Q(x,t,u)$ $=r(x,u)\,\delta t+e^{-\rho\delta t}V\bigl(x(t+\delta t),t+\delta t\bigr)$
$=r(x,u)\,\delta t+e^{-\rho\delta t}\mathbb{E}_{u^{\prime}\sim\pi(\cdot\mid x^{\prime})}\left[Q(x^{\prime},t+\delta t,u^{\prime})\right],$
@@
$\mathbb{E}_{u\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x,t,u)\right]\geq\mathbb{E}_{u\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x,t,u)\right].$
@@
$\begin{aligned} Q^{\pi_{\text{old}}}&=r_{0}\,\delta t+e^{-\rho\delta t}(\mathbb{E}_{u_{1}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{1},t_{1},u_{1})\right])\\ &\leq r_{0}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{1}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{1},t_{1},u_{1})\right]\right)\\ &=r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{2},t_{2},u_{2})\right]\right)\\ &\leq r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{2},t_{2},u_{2})\right]\right)\\ &=r_{0}\,\delta t+e^{-\rho t_{1}}r_{1}\,\delta t+e^{-\rho t_{2}}r_{2}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{3}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{3},t_{3},u_{3})\right]\right)\\ &\vdots\\ &\leq\int_{0}^{T}e^{-\rho t}r_{t}dt\\ &=Q^{\pi_{\text{new}}}.\end{aligned}.$
@@ A.4 Derivation of Definition 2
We consider the value function defined in Eq. (2) and follow the Proof of Lemma 3.1 to write out the dynamic programming principle of $V(x,t)$ as:
$V(x_{t},t)=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V(x_{t+\Delta t},t+\Delta t).$
where $u$ is the optimal control input. Taking the gradient with respect to $x$ on both sides using the chain rule:
$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x}f(x_{t},u_{t})^{\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$
@@
From VGI in definition 2, with an Euler step gives $x_{t+\Delta t}=x_{t}+f(x_{t},u_{t})\Delta t+o(\Delta t)$, we have the first-order-accurate VGI approximation:
$\nabla_{x}V(x_{t},t)=\nabla_{x}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\bigl[I+\nabla_{x}f(x_{t},u_{t})\Delta t\bigr]^{\!\top}\nabla_{x}V(x_{t+\Delta t},t+\Delta t).$
Rewrite VGI approximation as the affine map $\zeta=G(\zeta)$ with
$b=\nabla_{x}r(x_{t},u_{t})\Delta t,\;A=e^{-\rho\Delta t}\bigl[I+\nabla_{x}f(x_{t},u_{t})\Delta t\bigr]^{\!\top},\;\zeta=\nabla_{x}V(\cdot,\cdot).$
Assume the dynamics have a bounded Jacobian $\|\nabla_{x}f(x_{t},u_{t})\|\leq L_{f}$. Then
@@ A.6 Training Algorithm
12:   append $(x,t,u,r,x^{\prime})$ to local rollout $\mathcal{R}$
@@
18:  compute all losses $\mathcal{L}_{\mathrm{res}},\mathcal{L}_{\mathrm{term}},\mathcal{L}_{\mathrm{anchor}},\mathcal{L}_{\mathrm{vgi}}$ by Equation (5), (17), (13).
@@
22:   compute $A(x,t,u)$ for all $(x,t,u)\in\mathcal{R}$
23:   $\omega_{i}\leftarrow\omega_{i}-\alpha_{\pi}\,\nabla_{\omega_{i}}\bigl(-\mathbb{E}_{(x,t,u)\in\mathcal{R}}[\,A(x,t,u)\,\log\pi_{\omega_{i}}(u_{i}\mid x,t)\,]\bigr)$ by Equation (9).
@@ B.2 Continuous-Time MPE
We build on the standard MPE of [25], which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions. In the original MPE each control step advances the physics by a fixed time‐step $\Delta t_{\rm fixed}=0.1\,$s. To evaluate our continuous‐time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
@@ C.1 Value Gradient Comparison
Appendix D Hyper-parameters
@@
Value Network State + Time ($d+1$) FC(128) → FC(128) → FC(1), ReLU or Tanh
Dynamics Network State + Joint Action + Time ($d+na+1$) FC(128) → FC(128) → FC($d$), ReLU
Reward Network State + Joint Action + Time ($d+na+1$) FC(128) → FC(128) → FC(1), ReLU
PolicyNet Observation + Time + Time Interval ($o+2$) FC(128) → FC(128) → FC(64) → FC($a$), ReLU
Table 2 summarizes the architectures of the neural networks used in our VIP framework. All networks are implemented as FC layers with hidden size 128 unless otherwise specified. The value network takes the concatenation of the global state and time $(d+1)$ as input and outputs a scalar value. In our implementation, we use the Tanh activation function for the value network, as it provides smoother and more stable gradient propagation, which is critical for PINN-based value approximation. To validate this choice, we conducted an ablation study comparing Tanh and ReLU activations in the previous section.

### v3

@@ Appendix A Mathematical Proof
Given all $x\in\mathcal{X}$ with a small horizon $\Delta t>0$, we apply a Taylor expansion to the definition of $V(x)$ in Eq. 2 to derive the HJB PDE as follows:
$V(x)$ $=\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(\tau-t)}r(x,u)d\tau$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(s+\Delta t-t)}r(x(s+\Delta t),u(s+\Delta t))ds$
$=\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}\int_{t}^{\infty}e^{-\rho(s-t)}r(x(s+\Delta t),u(s+\Delta t))ds$
$\approx\max_{u\in\mathcal{U}}\int_{t}^{t+\Delta t}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}\int_{t+\Delta t}^{\infty}e^{-\rho(\tau-t)}r(x(\tau),u(\tau))d\tau$
$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+e^{-\rho\Delta t}\max_{u\in\mathcal{U}}V(x(t+\Delta t))$
$=\max_{u\in\mathcal{U}}r(x,u)\Delta t+(1-\rho\Delta t+o(\Delta t))\max_{u\in\mathcal{U}}(V(x)+\nabla_{x}V(x)^{\!\top}\!f(x,u)\Delta t+o(\Delta t))$
By canceling out $V(x)$ on both sides of the above equality, we obtain that
$-\rho V(x)\Delta t+\max_{u\in\mathcal{U}}(\nabla_{x}V(x)^{\!\top}\!f(x,u)+r(x,u))\Delta t+o(\Delta t)=0$
@@
$-\rho V(x)+\max_{u\in\mathcal{U}}(\nabla_{x}V(x)^{\!\top}\!f(x,u)+r(x,u))=0.$
Therefore, the $V(x)$ is the optimal solution to the following HJB PDEs:
$\rho V(x)+\nabla_{x}V(x)^{\!\top}\!f(x,u^{*})+r(x,u^{*})=0,$
Here optimal control input is $u^{*}=\operatorname*{arg\,max}_{u\in\mathcal{U}}\mathcal{H}(x,\nabla_{x}V(x))$, where $\mathcal{H}$ is the Hamiltonian defined as $\mathcal{H}=\nabla_{x}V(x)^{\!\top}\!f(x,u)+r(x,u)$. ∎
@@
$Q(x_{t},u_{t})=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}\,V\bigl(x_{t+\delta t}\bigr).$
For small $\delta t$ we have the first-order Taylor expansion in state:
$V\bigl(x_{t+\delta t}\bigr)=V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}\!f(x_{t},u_{t})\,\delta t+o(\delta t).$
@@
$Q(x_{t},u_{t})$ $=r(x_{t},u_{t})\,\delta t+\bigl(1-\rho\,\delta t+o(\delta t)\bigr)\Bigl[V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})\,\delta t+o(\delta t)\Bigr]$
$=r(x_{t},u_{t})\,\delta t+V(x_{t})+\bigl[\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})-\rho V(x_{t})\bigr]\delta t+o(\delta t).$
Subtract $V(x_{t})$ and discard the higher-order term:
$Q(x_{t},u_{t})-V(x_{t})=[-\rho V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t})]\delta t+o(\delta t).$
@@
$A(x_{t},u_{t})=\lim_{\delta t\to 0}\frac{Q(x_{t},u_{t})-V(x_{t})}{\delta t}=-\rho V(x_{t})+\nabla_{x_{t}}V(x_{t})^{\!\top}f(x_{t},u_{t})+r(x_{t},u_{t}).$
@@
$\pi_{\text{new}}(x_{t})=\arg\max_{u\in\mathcal{U}}Q^{\pi_{\text{old}}}(x_{t},u_{t}),$
@@
$Q(x_{t},u_{t})$ $=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}V\bigl(x_{t+\delta t}\bigr)$
$=r(x_{t},u_{t})\,\delta t+e^{-\rho\delta t}\mathbb{E}_{u^{\prime}\sim\pi(\cdot\mid x^{\prime})}\left[Q(x_{t+\delta t}^{\prime},u_{t+\delta t}^{\prime})\right],$
@@
$\mathbb{E}_{u\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{t},u_{t})\right]\geq\mathbb{E}_{u\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t},u_{t})\right].$
@@
$\begin{aligned} Q^{\pi_{\text{old}}}&=r_{t_{0}}\,\delta t+e^{-\rho\delta t}(\mathbb{E}_{u_{t_{1}}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t_{1}},u_{t_{1}})\right])\\ &\leq r_{t_{0}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{t_{1}}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{t_{1}},u_{t_{1}})\right]\right)\\ &=r_{t_{0}}\,\delta t+e^{-\rho t_{1}}r_{t_{1}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{2}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t_{2}},u_{t_{2}})\right]\right)\\ &\leq r_{t_{0}}\,\delta t+e^{-\rho t_{1}}r_{t_{1}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{t_{2}}\sim\pi_{\text{new}}}\left[Q^{\pi_{\text{old}}}(x_{t_{2}},u_{t_{2}})\right]\right)\\ &=r_{t_{0}}\,\delta t+e^{-\rho t_{1}}r_{t_{1}}\,\delta t+e^{-\rho t_{2}}r_{t_{2}}\,\delta t+e^{-\rho\delta t}\left(\mathbb{E}_{u_{t_{3}}\sim\pi_{\text{old}}}\left[Q^{\pi_{\text{old}}}(x_{t_{3}},u_{t_{3}})\right]\right)\\ &\vdots\\ &\leq\int_{t}^{\infty}e^{-\rho t}r_{t}dt\\ &=Q^{\pi_{\text{new}}}.\end{aligned}.$
@@ A.4 Derivation of Definition 2
We consider the value function defined in Eq. 2 and follow the Proof of Lemma 3.1 to write out the dynamic programming principle of $V(x_{t})$ as:
$V(x_{t})=r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}V(x_{t+\Delta t}).$
where $u_{t}$ is the optimal control input. Taking the gradient with respect to $x$ on both sides using the chain rule:
$\nabla_{x_{t}}V(x_{t})=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\nabla_{x_{t}}f(x_{t},u_{t})^{\top}\nabla_{x_{t+\Delta t}}V(x_{t+\Delta t}).$
@@
From the definition of VGI in Definition 2, applying a first-order Euler step gives
$x_{t+\Delta t}=x_{t}+f(x_{t},u_{t})\Delta t+o(\Delta t).$
This yields the first-order VGI approximation
$\nabla_{x_{t}}V(x_{t})=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t+e^{-\rho\Delta t}\bigl[I+\nabla_{x_{t}}f(x_{t},u_{t})\Delta t\bigr]^{\!\top}\nabla_{x_{t+\Delta t}}V(x_{t+\Delta t}).$
We can rewrite this approximation as an affine map
$\zeta=G(\zeta)$
where
$b=\nabla_{x_{t}}r(x_{t},u_{t})\Delta t,\;A=e^{-\rho\Delta t}\bigl[I+\nabla_{x_{t}}f(x_{t},u_{t})\Delta t\bigr]^{\!\top},\;\zeta=\nabla_{x}V(x).$
In Section 3.1, the joint dynamics function $f(x,u)$ is assumed time-invariant, which implies its Jacobian $\nabla_{x}f$ is also time-independent. Consequently, the matrix $A$ is time-invariant, and the update map can be written as a single contraction map
$G(\zeta)=b+\mathcal{A}\zeta,$
rather than a family of time-indexed maps $G_{t}$}. Assuming the dynamics have a bounded Jacobian,
$\|\nabla_{x}f(x_{t},u_{t})\|\leq L_{f},$
we obtain the bound
@@ A.6 Training Algorithm
12:   append $(x,u,r,x^{\prime})$ to local rollout $\mathcal{R}$
@@
18:  compute all losses $\mathcal{L}_{\mathrm{res}},\mathcal{L}_{\mathrm{anchor}},\mathcal{L}_{\mathrm{vgi}}$ by Eq. 17.
@@
22:   compute $A(x,u)$ for all $(x,u)\in\mathcal{R}$
23:   $\omega_{i}\leftarrow\omega_{i}-\alpha_{\pi}\,\nabla_{\omega_{i}}\bigl(-\mathbb{E}_{(x,u)\in\mathcal{R}}[\,A(x,u)\,\log\pi_{\omega_{i}}(u_{i}\mid x)\,]\bigr)$ by Eq. 8.
@@ B.2 Continuous-Time MPE
We build on the standard MPE of Lowe et al. (2017b), which simulates $N$ holonomic agents in a 2D world with simple pairwise interactions. In the original MPE each control step advances the physics by a fixed time‐step $\Delta t_{\rm fixed}=0.1\,$s. To evaluate our continuous‐time framework under irregular sampling, we modify the simulator so that at each step the integration interval is drawn randomly,
@@ C.1 Value Gradient Comparison
C.3 Stochastic Noise &\& Model Mismatch
Figure 12: Performance under Different Noise Levels.
Figure 12 evaluates the robustness of VIP under different levels of stochastic perturbation injected directly into the dynamics, according to
$x_{t+\Delta t}\;=\;f(x_{t},u_{t})\,\Delta t\;+\;\varepsilon_{t},\quad\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}I),$
where we consider three noise scales: $\sigma^{2}=0.1$ (Low), $0.5$ (Mid), and $1.0$ (High). With no or low noise, VIP maintains strong performance across both Predator–Prey and Cooperative Navigation, indicating that the learned continuous-time value function and its gradients remain stable under mild stochasticity. As the noise level increases, performance degrades noticeably, which is expected: our method is designed for deterministic continuous-time systems, and stochastic settings fundamentally alter the associated HJB equations (leading to stochastic HJB or HJB–Bellman–Fokker–Planck forms). Since VIP does not incorporate additional components for modeling diffusion terms or uncertainty, the degradation at high noise levels is consistent with the theoretical scope of our formulation.
C.4 Sensitivity Check for $\Delta t$
Figure 13: Average Return under Different $\Delta t$ in Coupled Oscillator .
[…]
@@ Appendix D Hyper-parameters
Value Network State ($d$) FC(128) → FC(128) → FC(1), ReLU or Tanh
Dynamics Network State + Joint Action ($d+na$) FC(128) → FC(128) → FC($d$), ReLU
Reward Network State + Joint Action ($d+na$) FC(128) → FC(128) → FC(1), ReLU
PolicyNet Observation + Time Interval ($o+1$) FC(128) → FC(128) → FC(64) → FC($a$), ReLU
Table 2 summarizes the architectures of the neural networks used in our VIP framework. All networks are implemented as FC layers with hidden size 128 unless otherwise specified. The value network takes the concatenation of the global state and time $(d)$ as input and outputs a scalar value. In our implementation, we use the Tanh activation function for the value network, as it provides smoother and more stable gradient propagation, which is critical for PINN-based value approximation. To validate this choice, we conducted an ablation study comparing Tanh and ReLU activations in the previous section.
Appendix E The Use of Large Language Models (LLMs)
We use LLMs as a writing assistant to polish/revise the paper.